OpenStudy (anonymous):

evaluate if a=2 and b=-5 3a^2

OpenStudy (anonymous):

I did that wrong.

OpenStudy (anonymous):

3 * 2 = 6 6 ^ 2 = 6 * 6 36.

OpenStudy (anonymous):

36

OpenStudy (anonymous):

you gotta make the a^2 first so it is 12

OpenStudy (anonymous):

O>O...No. No. 2^2 = 4....Oh yeah it is 12. o.o 2^2 = 4 3*4 = 12 12. Lol. I did not see that. I thought there were parenthesis.

OpenStudy (anonymous):

no problem